### v1

@@ 3 MPECopt: A globally convergent method for computing B-stationary points of MPECs
This section presents the MPEC Optimizer (MPECopt), an easy-to-implement active set method that solves a finite sequence of LPECs and BNLPs, and returns either a B-stationary point or a certificate of local infeasibility. Before providing all the details, we summarize the main ideas behind the new algorithm in Section 3.1. Thereafter, Section 3.2 discusses different formulations of LPECs, their properties, and how LPECs can be solved numerically. MPECopt consists of two phases. The first phase, described in Section 3.3, identifies a first feasible BNLP or detects local infeasibility. The second phase, described in Section 3.4, solves a sequence of BNLPs and LPECs and returns a B-stationary point.
@@ 3.1 Summary of the approach
Algorithm 1 provides a simplified pseudo-code of MPECopt. The algorithm starts from a feasible point $x^{0}\in\Omega$ and the corresponding ${\mathrm{BNLP}}({\mathcal{I}}_{1}^{0},{\mathcal{I}}_{2}^{0})$. If such a point is not immediately available, a Phase I algorithm can find one or certify that the MPEC is (locally) infeasible. We will present two different Phase I algorithms below. Given a point $x^{k}$, the algorithm requires solving the corresponding BNLP, compactly denoted by $\mathrm{BNLP}({\mathcal{I}}_{1}^{k},{\mathcal{I}}_{2}^{k})$, where ${\mathcal{I}}_{1}^{k}$ and ${\mathcal{I}}_{2}^{k}$ are the index sets defined in (7) at the point $x^{k}$. Moreover, LPECs similar to (6) with an additional trust region constraint $\|d\|_{\infty}\leq\rho$, $\rho>0$ are solved. The LPECs are compactly denoted by $\mathrm{LPEC}(x^{k},\rho)$, and are defined in more detail in the next section.
@@
[…]
3:Solve $\textrm{LPEC}(x^{k},\rho^{k,l})$ to obtain $d^{k,l}$
4:Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$ to obtain $x^{k,l}$If $f(x^{k,l})<f(x^{k})$
[…]
Phase II is summarized in lines 2 to 10 of Algorithm 1. It has major (outer) iterations that accept a new iteration $x^{k+1}$, corresponding to the solution of ${\mathrm{BNLP}}({\mathcal{I}}_{1}^{k,l},{\mathcal{I}}_{2}^{k,l})$, only if it has a better objective than the solution of ${\mathrm{BNLP}}({\mathcal{I}}_{1}^{k},{\mathcal{I}}_{2}^{k})$. The minor (inner) iterations try to find a BNLP whose solution improves to objective value. The solution of the LPEC is used to provide an active set guess for the BNLP, i.e., after an LPEC solve we solve ${\mathrm{BNLP}}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$, where $d^{k,l}$ is solution of the current LPEC and $x^{k}$ is the current iterate. If a step is rejected then $\mathrm{LPEC}(x^{k},\rho^{k,l+1})$ is re-solved in the next inner iteration but now with a smaller trust region radius $\rho^{k,l+1}$. This eventually provides a better active set guess for the next BNLP.
It is important to note that the LPEC steps $x^{k}+d^{k,l}$ are never used to update the iterate because this might be infeasible. The LPEC is only used as an oracle for obtaining a new active set guess and to determine when to terminate. This approach has several advantages. Starting from a feasible point $x^{k}$, the LPEC steps $x^{k}+d^{k,l}$ can both increase the actual objective $f(x^{k}+d^{k,l})$ and make the new point infeasible, i.e., $x^{k}+d^{k,l}\notin\Omega$. This would require a sophisticated globalization strategy to ensure sufficient decrease and that $x^{k}+d^{k,l}$ is sufficiently feasible. Fortunately, this can be avoided by solving only BNLPs with a globalized NLP solver, which starts from a feasible point $x^{k}$, returns a new feasible point with a better objective value, or verifies the stationarity of $x^{k}$ for the current BNLP.
3.2 The LPEC suproblems
Next, we discuss and compare different formulations for the LPEC used in our method, and then we discuss how they are efficiently solved in practice.
@@ 3.2.1 On different LPEC formulations
Given a feasible point $x^{k}$, Algorithm 1 requires frequently solving the following $\mathrm{LPEC}(x^{k},\rho)$:
@@
s.t. $c(x^{k})+\nabla c(x^{k})^{\top}d\geq 0,$ (12b)
$x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{0+}(x),$ (12c)
$x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x),$ (12d)
@@
This LPEC is similar to the LPEC (6) in Def. 2.1 with the following additions. Additional to (6c) and (6d), here we add the constraint $x_{2,i}^{k}+d_{2,i}\geq 0$ and $x_{1,i}^{k}+d_{1,i}\geq 0$ in (12c) and (12d), respectively. This ensures that for any solution $d$ of (12) the point $x^{k}+d$ satisfies the complementarity constraints, which is necessary for a meaningful definition of the active sets in  (7). We also add a trust region constraint for the step $d$ in (12f), where $\rho>0$ is the trust region radius. This constraint ensures that the LPEC is always bounded. We use an $\ell_{\infty}$ norm for the step since it results in simple lower and upper bounds on the step.
If $d=0$ is a local minimizer of (12), then $x^{k}$ is feasible for the MPEC (1) and B-stationary. All additional constraints present in (12) but not in  (6) are inactive at $d=0$. Thus, the feasible sets of the two LPECs coincide locally.
If $x^{k}$ is feasible but not B-stationary (i.e., $d=0$ is not a local minimizer of (12)), then $\nabla f(x^{k})^{\top}d<0$ and the LPEC solution computes a descent direction. In this case, we take the sets ${\mathcal{I}}_{1}(x^{k}+d)$ and ${\mathcal{I}}_{2}(x^{k}+d)$ as the new active set guess for the next BNLP.
Note that  (12e) considers only complementarity constraints for $i\in{\mathcal{I}}_{00}(x^{k})$. If the cardinality of this index set is small, the resulting LPEC is usually not hard to solve. However, if a point $x^{k}$ is infeasible – which often happens in Phase I – then it is unclear how to best define meaningful index sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+}$, ${\mathcal{I}}_{00}$ and the corresponding LPEC. While some active set identification strategies from infeasible points exist, they require strong assumptions to reliably identify meaningful active sets Lin2006 . Alternatively, some authors consider all constraints Kirches2022 ; Leyffer2007 , which leads to the following LPEC:
s.t. $c(x^{k})+\nabla c(x^{k})^{\top}d\geq 0,$ (13b)
$\|d\|_{\infty}\leq\rho.$ (13d)
We call the LPEC (12) the “reduced” LPEC and (13) the “full” LPEC. The latter contains the feasible set of the former and does not require active set estimates for an infeasible $x^{k}$. Moreover, at infeasible points $x^{k}$ the objective function value of (13) can be positive, and by reducing the trust region radius $\rho$ it can become nonnegative, but not the other way around. At infeasible points $x^{k}$, the trust region radius $\rho$ must be large enough to be compatible with the complementarity constraints (13c). One disadvantage of (13) over (12) is that has potentially many complementarity constraints that must be handled, making it possibly harder to solve. However, unlike the reduced version, the full LPEC can easily be used in both Phases I and II. The solution sets of the two LPECs are related as follows.
Lemma 3.1
$0<\rho<\bar{\rho}=\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\},$ (14)
the sets of the local minimizers of  (12) and (13) are identical. In the special case of ${\mathcal{I}}_{00}=\{1,\ldots,m\}$ the reduced and full LPEC coincide and in that case, $\bar{\rho}=\infty$.
Proof. Denote by $\Omega^{\mathrm{R}}$ the feasible set of the LPEC (12). Note that the feasible set of  (13) is
$\begin{split}\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}&\bigcup_{i\in{\mathcal{I}}_{0+}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\\ &\bigcup_{i\in{\mathcal{I}}_{+0}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}.\end{split}$ (15)
Now observe that, for example, for any $i\in{\mathcal{I}}_{+0}^{k}$ it holds that $x_{1,i}^{k}>0$ and $x_{2,i}=0$. If $\rho$ is chosen such that $d_{1,i}<\rho<x_{1,i}^{k}$ then $x_{1,i}^{k}+d_{1,i}>0$ and $0\leq x_{1,i}^{k}+d_{1,i}\perp d_{2,i}\geq 0$ can be only satisfied for $d_{2,{i}}=0$. Thus, the branch $\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$ is infeasible. Similar reasoning holds for any $i\in{\mathcal{I}}_{0,+}^{k}$. Therefore, by picking $\rho<\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\}$, none of the terms in the unions in (15) is feasible, hence $\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}$. As the two LPECs have the same objective, their sets of local minimizers coincide. ∎
In simpler words, a sufficiently small trust region radius cuts off all branches that are not biactive at $x^{k}$, and the two LPECs have the same feasible set. Therefore, if some $x^{*}\in\Omega$ is a local minimizer of the MPEC (1), then $d=0$ is a local minimizer of (13).
@@
Next, we show that the assumption of MPEC-Abadie CQ, i.e., that the tangent cone ${\mathcal{T}}_{\Omega}(x)$ is equal to MPEC-linearized feasible cone $\mathcal{F}_{\Omega}(x)$, cannot be relaxed if one wishes to verify B-stationarity by solving an LPEC. Recall that the MPEC-Abadie CQ is implied by the MPEC-MFCQ Flegel2005a .
@@
Figure 1: Illustration of Example 3.1. If a constraint qualification does not hold, the LPEC cannot verify B-stationary.
@@
$\underset{x\in{\mathbb{R}}^{2}}{\mathrm{min}}\;\quad$ $-2x_{1}+x_{2}$
s.t. $x_{2}-x_{1}^{2}\geq 0$
$0\leq x_{1}\perp x_{2}\geq 0$
The point $x^{*}=(0,0)$ is the global optimum of this problem and thus B-stationary. It can be seen that the tangent cone of this problem is ${\mathcal{T}}_{\Omega}(x^{*})=\{d\in{\mathbb{R}}^{2}\mid d_{1}=0,d_{2}\geq 0\}$. The corresponding LPEC reads as:
@@
Clearly, ${\mathcal{T}}_{\Omega}(x^{*})\subset\mathcal{F}_{\Omega}(x^{*})$, which means that the MPEC-Abadie CQ does not hold at $x^{*}$. The only solution of (17) is $d=(\rho,0)$. Therefore, this LPEC cannot verify B-stationarity for any $\rho>0$. Fig. 1 illustrates the feasible sets of the MPEC and the LPEC for different $\rho$.
Next, we directly compare the two LPEC formulations and highlight some of their strengths and weaknesses in two examples.
(a) $\rho=1.1$.
(b) $\rho=0.6$.
(c) $\rho=0.2$.
Figure 2: Illustration of the MPEC  (18). The yellow stars denote the two stationary points. The figure illustrates the LPEC’s (13) trust region, objective gradient, and optimal solution $d$ for different values of $\rho$.
Example 3.2 (Verifying B-stationarity)
Consider the following MPEC:
$\min_{x\in{\mathbb{R}}^{2}}\ (ax_{1}-1)^{2}+(x_{2}-1)^{2}\quad\textnormal{s.t.}\quad 0\leq x_{1}\perp x_{2}\geq 0,$ (18)
with $a>0$. This problem has two B-stationary points $\bar{x}=(\frac{1}{a},0)$ and $\hat{x}=(0,1)$, with $f(\bar{x})=f(\hat{x})=1$. We want to verify the B-stationarity of $\bar{x}$ with an LPEC solver that returns a globally optimal solution (which we often do in practice, cf. Section 3.2.2). Figure 2 illustrates this situation. For $\rho>\frac{1}{a}$, the solution of the full LPEC (13) will select the other branch, with the feasible set $x_{1}=0,x_{2}\geq 0$. However, the local optimum of this BNLP has the same objective value as the initial BNLP and the step will be rejected and the trust region radius reduced. The steps will be rejected until $\rho<\frac{1}{a}$, which for $a\ll 1$ may require many steps. On the other hand, the reduced LPEC (12) is just an LP and the B-stationary point is verified after the first LP is solved. For $\rho<\frac{1}{a}$, the full and reduced LPEC have the same feasible set, cf. Lemma 3.1.
A straightforward way to identify a B-stationary point in the example above is to exploit the fact that if a local optimizer of a BNLP $x^{*}$ has an empty biactive set ${\mathcal{I}}_{00}(x^{*})=\emptyset$, then this point is B-stationary. However, in practice, on a finite precision machine, the active sets must be determined with some threshold, e.g.,
${\mathcal{I}}_{00}^{\epsilon}(x)=\{i\mid x_{1,i}\leq\epsilon,x_{2,i}\leq\epsilon\}.$ (19c)
If in (18), one has $\epsilon>\frac{1}{a}$, a wrong conclusion about stationarity will be made, whereas the full LPEC does not need active set identification and always correctly identifies the B-stationary point.
@@
This problem has two B-stationary points $\bar{x}=(1,0)$ and $\hat{x}=(0,1)$ with the respective objective values $f(\bar{x})=1$ and $f(\hat{x})=4$. In Fig. 3a, if $\rho$ is sufficiently small or the reduced LPEC is used, the point $\hat{x}$ with $f(\hat{x})=4$ is verified as B-stationary point. On the other hand, with a larger trust region radius, as depicted in Fig. 3b, the globally optimal solution of the full LPEC identifies a BNLP with a lower objective value $f(\bar{x})=1$. Conversely, if we start at $\bar{x}$, the full LPEC can identify a BNLP with a larger objective, as depicted in Fig. 3c. However, this step will be rejected, and the trust region radius reduced until $\bar{x}$ is verified as B-stationary.
We make several observations from this example. First, solving the full LPEC to global optimality can help to find an objective with a lower value. Of course, as illustrated in Fig. 3c, there is no guarantee that this will always be the case. Second, if the trust region is not small enough, there is no guarantee that the descent direction corresponding to the global optimum of the LPEC leads to a BNLP with a lower objective value. Third, observe that when starting from $\hat{x}$ and then solving the BNLP predicted by the full LPEC as in Fig. 3b, the point $\hat{x}$ is not a feasible initial guess for the new BNLP, which may lead to step rejection if an NLP solver is not able to solve the new BNLP.
Next, we give sufficient conditions under which an ${\mathrm{LPEC}}(x^{k},\rho)$ predicts a BNLP for which $x^{k}$ is always a feasible point.
@@
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1) and suppose that the MPEC-Abadi CQ holds at this point. Let $d$ be any feasible point of the $\mathrm{LPEC}(x^{k},\rho)$ in (13) for $\rho\in(0,\bar{\rho})$, where $\bar{\rho}$ is defined in (14). Then $x^{k}$ is feasible in $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d),{\mathcal{I}}_{2}(x^{k}+d))$.
Proof. It follows from Lemma 3.1 that for $\rho\in(0,\bar{\rho})$ both the full and reduced LPEC have the same feasible set. The relation (10) is also true for the respective tangent cones of the MPEC and auxiliary NLPs. Moreover, due to the MPEC-Abadi CQ, the relation (10) also holds for the MPEC-linearized feasible cones, i.e., the feasible set of the ${\mathrm{LPEC}}(x^{k},\rho)$. Thus, the proof follows directly from (10). ∎
Note that, since only feasible points $d$ are needed in Lemma 3.2, the assertion is true if we replace the cost gradient $\nabla f(x^{k})$ in $\mathrm{LPEC}(x^{k},\rho)$ with any vector $a\in{\mathbb{R}}^{n}$. Lemma 3.2 will be of great importance in the convergence analysis of MPECopt done in Section 4.
In summary, the full LPEC is easy to formulate at both feasible and infeasible points. The reduced LPEC is more suitable for feasible points and may be easier to solve. The full LPEC can find a better local optimizer at the cost of starting the next BNLP from an infeasible point. For a sufficiently small trust region radius, the feasible sets of the two LPECs coincide. Moreover, any feasible point of a reduced LPEC leads to an active set guess that corresponds to a feasible BNLP. In conclusion, in a practical implementation, one should use the full LPEC with a reasonably small trust region radius.

### v2

@@ 3 MPECopt: A globally convergent method for computing B-stationary points of MPECs
This section presents the MPEC Optimizer (MPECopt), an easy-to-implement active set method that solves a finite sequence of LPECs or LPs, and BNLPs. It returns either a B-stationary point or a stationary point of the problem of minimizing the constraint infeasibility. The latter is a common outcome of nonlinear programming algorithms Fletcher2002 ; Waechter2006 .
Before providing all the details, we summarize the main ideas behind the new algorithm in Section 3.1. Thereafter, Section 3.2 discusses different formulations of LPECs, their properties, and how LPECs can be solved numerically. MPECopt consists of two phases. Details of Phase I and Phase II are discussed in Sections 3.3 and  3.4, respectively.
@@ 3.1 Summary of the approach
1: Call Phase I to find a feasible point $x^{0}$ of the MPEC (1)
2: while $d^{k,l}\neq 0$ do $\triangleright$ major iterations
3:   $x^{k,0}=x^{k}$
4:   for $l=0,\ldots,$ do $\triangleright$ minor iterations
5:    Solve $\textrm{LPEC}(x^{k},\rho^{k,l})$ in (15) to obtain $d^{k,l}$
6:    if $d^{k,l}=0$ then terminate $\triangleright$ B-stat. point found
7:    Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$ to obtain $x^{k,l}$
8:    if BNLP unbounded then terminate $\triangleright$ MPEC is unbounded
9:    else if $f(x^{k,l})<f(x^{k})$ then
10:      Set $x^{k+1}\!=\!x^{k,l}$, k=k+1k\!=\!k\!+\!1, increase trust-region radius $\rho^{k+1,0}$, and break for-loop
11:    else
12:      Reduce trust-region radius $\rho^{k,l+1}$
Algorithm 1 provides simplified pseudo-code for MPECopt, which consists of two phases. Phase I either finds an MPEC-feasible point $x^{0}\in\Omega$ or a stationary point of the problem of minimizing the constraint infeasibility. Phase II is summarized in lines 2–11 of Algorithm 1, while the details of our actual implementation are given in Section 3.4. The outcome of Phase II is that it finds a B-stationary point of the MPEC or detects that the problem is unbounded.
If a feasible point is found in Phase I, Phase II creates a finite sequence of iterates $x^{k}\in\Omega$ by solving BNLPs and LPECs. If Phase II finds that any BNLP is unbounded, then the MPEC is unbounded, and the algorithm terminates. The LPECs are similar to (6), but they are defined at the linearization point $x^{k}$ and include an additional trust-region constraint $\|d\|_{\infty}\leq\rho$, $\rho>0$. These LPECs are compactly denoted by $\mathrm{LPEC}(x,\rho)$; see Section 3.2 for more details.
At each major iteration $k$, an inner loop with counter $l$ solves $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$ until either a feasible point $x^{k+1}=x^{k,l}$ with an improved objective value is found (lines 7–10), or the algorithm terminates due to $d^{k,l}=0$ (line 6), or unboundedness (line 8).
@@
Note that the LPEC steps $x^{k}+d^{k,l}$ are never used to update iterates directly. This avoids cases where, under nonlinear constraints, $x^{k}+d^{k,l}\notin\Omega$ while $f(x^{k}+d^{k,l})$ increases, which would require sophisticated globalization for the LPEC steps. Instead, we accept only BNLP stationary points $x^{k,l}$, which remain feasible for the MPEC (1).
3.2 The LPEC subproblems
In Section 3.2.1 we discuss and compare two different formulations for the LPEC used in our method, and in Section 3.2.2 we discuss how they are efficiently solved.
@@ 3.2.1 On different LPEC formulations
The first formulation in inspired by Def. 2.1. Given a feasible point $x^{k}\in\Omega$, Phase II in Algorithm 1 requires solving the following $\mathrm{LPEC}(x^{k},\rho)$:
@@
s.t. $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,$ ${\forall i\in{\mathcal{A}}(x^{k})}$ (12b)
$x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{0+}(x^{k}),$ (12c)
$x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{k}),$ (12d)
@@
Compared to (6) in Def. 2.1, the LPEC above has the following modifications. Additional to (6c) and (6d), here we include the constraint $x_{2,i}^{k}+d_{2,i}\geq 0$ and $x_{1,i}^{k}+d_{1,i}\geq 0$ in (12c) and (12d), respectively. This ensures that for any solution $d$ of (12) the point $x^{k}+d$ satisfies the complementarity constraints, which is necessary for a meaningful definition of the active sets in  (7). Moreover, we add a trust region constraint for the step $d$ in (12f), where $\rho>0$ is the trust region radius. This constraint guarantees that the LPEC is bounded. We use an $\ell_{\infty}$ norm for the step, yielding simple lower and upper bounds on $d$.
If $d=0$ is a global minimizer of (12), then none of the additionally introduced constraints is active, and moreover, if the MPEC-ACQ holds, then by Theorem 2.1, $x^{k}$ is a B-stationary point of MPEC (1). On the other hand, if $x^{k}$ is not B-stationary (i.e., $d=0$ is not a minimizer of (12)), and if ${\mathcal{A}}(x^{k}+d)={\mathcal{A}}(x^{k})$, then $\nabla f(x^{k})^{\top}d<0$ and $d$ is a first-order descent direction. In this case, we take the sets ${\mathcal{I}}_{1}(x^{k}+d)$ and ${\mathcal{I}}_{2}(x^{k}+d)$ as the new active set guess for the next BNLP to be solved in Phase II; more details on the actual selection of these index sets are provided below.
@@
Next, we show that the assumption of MPEC-ACQ (i.e., that ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$), cannot be relaxed if one wishes to verify geometric B-stationarity (cf. Theorem 2.1) by solving the LPEC 12, which verifies algebraic/linearized B-stationarity (cf. Def. 2.1).
@@
Figure 1: Illustration of Example 3.1. If a constraint qualification does not hold, the LPEC cannot verify geometric B-stationary.
@@
[…]
The point $x^{*}=(0,0)$ is the global optimum of this problem and thus geometric B-stationary. It can be seen that the tangent cone of this problem is ${\mathcal{T}}_{\Omega}(x^{*})=\{d\in{\mathbb{R}}^{2}\mid d_{1}=0,d_{2}\geq 0\}$. The corresponding LPEC reads as:
@@
Clearly, ${\mathcal{T}}_{\Omega}(x^{*})\subset\mathcal{F}_{\Omega}(x^{*})$, which means that the MPEC-ACQ does not hold at $x^{*}$. The only solution of (14) is $d=(\rho,0)$. Therefore, this LPEC cannot verify geometric B-stationarity for any $\rho>0$. Fig. 1 illustrates the feasible sets of the MPEC and the LPEC for different $\rho$.
The LPEC (12) is well-defined at feasible points. However, in the Phase I algorithm of Section 3.3, we must also consider LPECs at infeasible points, where defining the index sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+},{\mathcal{I}}_{00}$ is unclear. Existing strategies for identifying active sets from infeasible points often require overly restrictive assumptions Lin2006 . Inspired by (Scholtes2004, , Sec. 3.2), alternative approaches instead include all linearized constraints, yielding the LPEC Kirches2022 ; Leyffer2007 :
s.t. $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,$ $\forall i\in\{1,\ldots,n_{c}\},$ (15b)
$\|d\|_{\infty}\leq\rho.$ (15d)
We refer to the LPEC (12) as the “reduced” LPEC and (15) as the “full” LPEC. For feasible points $x^{k}$, the full LPEC contains the feasible set of the reduced one. Moreover, since $d=0$ is always feasible for $x^{k}\in\Omega$, the optimal objective values of both (12) and (15) are nonpositive, regardless of the trust region radius $\rho>0$. At infeasible points $x^{k}\notin\Omega$, the optimal objective value of (15) can take any value in $\mathbb{R}$}. Moreover, at such points, the trust-region radius $\rho$ must be large enough to ensure that the LPEC is feasible, cf. Theorem 4.1.
Let us introduce the following notation shorthand ${\mathcal{I}}_{+0}^{k}:={\mathcal{I}}_{+0}(x^{k})$ and ${\mathcal{I}}_{0+}^{k}:={\mathcal{I}}_{0+}(x^{k})$. Denote the set of inactive inequality constraints by $\bar{{\mathcal{A}}}^{k}:=\bar{{\mathcal{A}}}(x^{k})=\{i\in\{1,\ldots,n\}\mid c_{i}(x)>0\}$. For a feasible point $x^{k}\in\Omega$, the solution sets of the two LPECs are related as follows. The proof can be found in Appendix A.
Lemma 3.1
$\begin{split}0<\rho<\bar{\rho}=\min\Big\{&\{{x^{k}_{1,i}}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{{x^{k}_{2,i}}\mid i\in{\mathcal{I}}_{0+}^{k}\}\\ &\cup\{\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}\mid i\in\bar{{\mathcal{A}}}^{k},\|\nabla c_{i}(x^{k})\|\neq 0\}\Big\},\end{split}$ (16)
the sets of the local minimizers of  (12) and (15) are identical. In the special case of ${\mathcal{I}}_{00}^{k}=\{1,\ldots,m\}$ and $\bar{{\mathcal{A}}}^{k}=\emptyset$ the reduced and full LPEC coincide and in that case, $\bar{\rho}=\infty$.
In simpler terms, at a feasible point $x^{k}$ and for a sufficiently small trust-region radius, all branches that are not biactive at $x^{k}$ are eliminated, and the linearizations of inactive constraints remain inactive. Thus, the feasible sets of the two LPECs coincide. On the other hand, for $\rho\geq\bar{\rho}$, the global minimizer $d=0$ of (12) may correspond only to a local minimizer of (15), cf. Example 3.2. Hence, global optimality of $d=0$ is sufficient but not necessary to verify B-stationarity via (15), cf. Scholtes2004 .
[…]
@@
The example has two S-stationary points, $\bar{x}=(1,0)$ and $\hat{x}=(0,1)$, both with empty biactive sets, and respective objective values $f(\bar{x})=1$ and $f(\hat{x})=4$. In Fig. 2a, if $\rho$ is sufficiently small or the reduced LPEC is used, the point $\hat{x}$ with $f(\hat{x})=4$ is verified as B-stationary point. On the other hand, with a larger trust region radius, as depicted in Fig. 2b, the globally optimal solution of the full LPEC identifies a BNLP with a lower objective value $f(\bar{x})=1$. Conversely, if we start at $\bar{x}$, the full LPEC can identify a BNLP with a larger objective, as depicted in Fig. 2c. However, this step will be rejected, and the trust region radius reduced until $\bar{x}$ is verified as B-stationary.
[…]
@@
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1) and suppose that the MPEC-ACQ holds at this point. Let $d$ be any feasible point of the $\mathrm{LPEC}(x^{k},\rho)$ in (15) for $\rho\in(0,\bar{\rho})$, where $\bar{\rho}$ is defined in (16). Then $x^{k}$ is a feasible point of $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d),{\mathcal{I}}_{2}(x^{k}+d))$.
A proof is given in Appendix A.